OpenStudy (anonymous):

sec theta/sin theta - sin theta/cos theta I've gotten it down to 1-sin ^2 theta/ sin theta

OpenStudy (anonymous):

\[1-\sin ^{2}\theta/\sin \theta\]

OpenStudy (whpalmer4):

1-sin^x = cos^x, remember?

OpenStudy (whpalmer4):

Don't forget that \[\cot x =\frac{ \cos x}{\sin x} = \frac{1}{\tan x}\]

OpenStudy (anonymous):

\[ \frac{\sec(\theta)}{\sin(\theta)} - \frac{\sin(\theta)}{\cos(\theta)} \equiv \frac{1}{\sin(\theta)\cos(\theta)} - \frac{\sin(\theta)}{\cos(\theta)} \equiv \frac{1 - \sin^2(\theta)}{\sin(\theta)\cos(\theta)} \equiv \frac{\cos^2(\theta)}{\sin(\theta)\cos(\theta)} \\ \frac{\cos^2(\theta)}{\sin(\theta)\cos(\theta)} \equiv \frac{\cos(\theta)}{\sin(\theta)} \equiv \cot(\theta) \]

OpenStudy (anonymous):

I swear I stare straight at these things. i'm so sorry mr. palmer. :(

OpenStudy (anonymous):

or miss...idk o.o

OpenStudy (whpalmer4):

my friends call me Bill

OpenStudy (anonymous):

Natalie :) pleasure to meet you Bill.

OpenStudy (anonymous):

OKAY go rescue someone else, i'll go to the next one!